# Existence and Uniqueness for a Porous-Medium-Type Diffusion Model with Time-Dependent Growth

### v2

@@ 3. Derivation of the Model via Conservation Laws
By choosing the specific function $\varphi(v)=v^{\alpha}$, the equation (1.1) is obtained after rearrangement. The interpretations are as follows: the diffusive term $\Delta v^{\alpha}$ describes the dispersal of individuals in a habitat (representing the way in which the population spreads out in space), while the term $\mu\left(t\right)v$ represents the effect of population growth or mortality, modeling the birth and death processes. The boundary conditions (1.2) and the initial condition (1.3) complement the model by ensuring that the initial population distribution is predefined and that there is no flux of individuals across the domain’s boundary.
4. Conclusions
In this work, we have established a comprehensive theoretical and computational framework for the nonlinear diffusion equation of Gurtin–MacCamy type with time-dependent growth rates. Our main contributions can be summarized as follows:
@@ 4. Conclusions
    $v(x,t)=S(t)u^{1/\alpha}(x),$
    where the spatial profile $u(x)$ satisfies a sublinear elliptic equation and the temporal function $S(t)$ is governed by a Bernoulli-type ODE. This separation-of-variables approach provides explicit analytical expressions that are directly implementable.
@@
    Theorem 1.3: We extended the existence and uniqueness theory to the general non-separable case using the monotone iteration method based on subsolution–supersolution techniques. This result demonstrates that solutions remain bounded within the ordered interval $v_{-},v_{+}]$ for arbitrary initial data satisfying the compatibility conditions.
@@
    Theorem 1.2: We established a more general comparison principle, showing that if the initial population $v_{0}$ is smaller than $w_{0}$ and the growth rate $\mu_{1}(t)$ is smaller than $\mu_{2}(t)$, then the population level $v(x,t)$ remains bounded by $w(x,t)$ for all subsequent times.
@@
    Numerical Validation: Python implementations of both theorems verify the consistency between the analytical separable solution and the general iterative approach. The simulations demonstrate that the iterative scheme converges to the exact separable form when applicable.
The parameter sensitivity analysis reveals that: the nonlinear exponent $\alpha$ governs the diffusion rate and solution regularity; the time-dependent growth rate $\mu(t)$ controls the temporal dynamics and long-time behavior; and the initial parameter $\gamma$ determines the amplitude scaling of the solution.
The proposed framework has broad applicability beyond population dynamics, including thermal diffusion in non-homogeneous materials and the modeling of bacterial contamination in healthcare environments. Future research directions include: extension to higher-dimensional domains with complex geometries; incorporation of spatially varying diffusion coefficients; and analysis of equations with nonlinear boundary conditions.
Details of the full Python implementations are provided in Appendices AC, enabling reproducibility and practical application of the theoretical results.
Acknowledgments
Funding
This research received no external funding.
Data Availability
Data sharing is not applicable since the article is purely theoretical in nature.
Declarations
Conflicts of Interests The author declares no conflict of interest.
Appendix A Python Implementation of Theorem 1.1
$v(x,t)=S(t)\cdot u^{1/\alpha}(x),$
as established in Theorem 1.1, is publicly available at: https://github.com/coveidragos/PorousMediumEquation/blob/main/Theorem_First.py.
Appendix B Python Implementation of Theorem 1.3
The Python code available at: https://github.com/coveidragos/PorousMediumEquation/blob/main/Theorem_Second.py implements Theorem 1.3 using the subsolution–supersolution framework. In the special case where the solution admits a separable form
$v(x,t)=S(t)\cdot u^{1/\alpha}(x),$
Appendix C Python Implementation of Remark 1.4
The Python script available at: https://github.com/coveidragos/PorousMediumEquation/blob/main/Theorem_Second.py extends the implementation from Section B to illustrate Theorem 1.3 under the general initial condition proposed in Remark 1.4. In particular, the parabolic solver is initialized with a weighted average of the sub- and supersolutions,
$v_{0}(x)=\frac{7}{8}v_{-}(x,0)+\frac{1}{8}v_{+}(x,0),$
thereby demonstrating that the monotone iteration scheme robustly captures the existence and uniqueness of solutions for arbitrary initial data within the ordered interval $v_{-},v_{+}]$.
Appendix D Image Denoising Application and Comparison with Perona–Malik
From an application perspective, the nonlinear diffusion models studied in this work, particularly the Gurtin–MacCamy (GM) type equations, find significant utility in the field of digital image processing. Specifically, the power-law diffusion mechanism governed by (1.1) can be effectively employed for image denoising. While classical linear diffusion (the heat equation) tends to blur important structural features like edges, the nonlinear approach preserves sharp transitions while smoothing homogeneous regions.

### v3

@@ 4. Derivation of the Model via Conservation Laws
By choosing the specific function $\varphi(v)=v^{\alpha}$, the equation (1.1) can be obtained after rearrangement. The interpretations are as follows: the diffusive term $\Delta v^{\alpha}$ describes the dispersal of individuals in a habitat (representing the way in which the population spreads out in space), whereas the term $\mu\left(t\right)v$ represents the effect of population growth or mortality, modeling the birth and death processes. The boundary conditions (1.2) and the initial condition (1.3) complement the model by ensuring that the initial population distribution is predefined and that there is no flux of individuals across the boundary of the domain.
5. Cross-Disciplinary Relevance
The Bernoulli localization principle developed in this paper is not confined to a single application domain. In this section we show how Equation (1.1) and its structural decomposition via the Bernoulli ODE are relevant to physics, biology, economics, and engineering. For each discipline we identify the physical model, specify what the variables represent, explain how the localization principle applies, and state a concrete, verifiable prediction.
5.1. Physics: thermal diffusion in porous media
In the physics of porous media, the temperature $v(x,t)$ in a material with temperature-dependent thermal conductivity $k(v)=\alpha v^{\alpha-1}$ satisfies the nonlinear heat equation
$\frac{\partial v}{\partial t}=\nabla\cdot(k(v)\nabla v)+Q(t)\,v=\Delta v^{\alpha}+Q(t)\,v,$ (5.1)
where $Q(t)$ is a time-dependent volumetric heat source (or sink). This is precisely Equation (1.1) with $\mu(t)=Q(t)$. The power-law dependence $k(v)=\alpha v^{\alpha-1}$ arises in radiative heat transfer ($\alpha=4$, Stefan–Boltzmann regime) and in the thermal conductivity of certain polymers and ceramics [2, 12].
Verifiable prediction. For a porous medium with thermal conductivity $k(v)=\alpha v^{\alpha-1}$ (e.g., $\alpha=4$ for radiation-dominated transfer), initial temperature profile $v_{0}(x)$, and time-dependent heat source $Q(t)=Q_{0}/(1+t)$, the Bernoulli localization principle predicts that the temperature at any point $x$ and time $t$ satisfies
$S(t)\,u_{-}^{1/\alpha}(x)\leq v(x,t)\leq S(t)\,u_{+}^{1/\alpha}(x),$
where $u_{-}$ and $u_{+}$ are the elliptic sub- and supersolutions on the spatial domain, and $S(t)$ is computed from the closed-form expression (1.4) with $\mu(t)=Q_{0}/(1+t)$. These bounds can be tested against numerical simulations or experimental measurements of heat flow in porous ceramics.
5.2. Biology: population dynamics and tumor growth
Equation (1.1) was originally introduced by Gurtin and MacCamy [16] to model the dispersal of biological populations with density-dependent diffusion. In this context, $v(x,t)$ represents the population density, $\alpha>1$ captures the fact that organisms in crowded regions disperse more rapidly (the “slow diffusion” regime, where the diffusion coefficient vanishes when the density is zero, leading to compact support and finite-speed propagation), and $\mu(t)$ represents a time-varying net reproduction rate.
A particularly important biological application is tumor growth modeling. The density of tumor cells in a tissue region $\Omega$ evolves according to a porous-medium-type equation, where the power-law diffusion $\Delta v^{\alpha}$ captures the mechanical pressure exerted by the growing tumor mass, and $\mu(t)$ represents a time-varying proliferation rate that may decrease due to nutrient depletion or therapeutic intervention [18].
Verifiable prediction. For a tumor with initial density profile $v_{0}(x)$ and proliferation rate $\mu(t)=\mu_{0}e^{-\beta t}$ (representing exponential decay due to chemotherapy), the maximum tumor cell density at time $t$ is bounded above by $S(t)\max_{x}u_{+}^{1/\alpha}(x)$, where $S(t)$ is the explicit Bernoulli solution (1.4) with the given $\mu(t)$. This prediction can be tested against in vitro spheroid growth data.
5.3. Economics and social sciences: diffusion of innovations
The spread of a new technology, product, or idea through a population can be modeled by nonlinear diffusion equations [23]. In this framework, $v(x,t)$ represents the density of adopters at location $x$ and time $t$, the power-law diffusion $\Delta v^{\alpha}$ captures the nonlinear “peer-influence” effect (adoption accelerates when adopter density is high), and $\mu(t)$ represents a time-varying exogenous stimulus (e.g., advertising, policy incentives, or media coverage).
The domain $\Omega$ represents a geographical region, and the Dirichlet boundary condition $v=0$ on $\partial\Omega$ models the fact that adoption does not penetrate beyond the market boundary. The Bernoulli localization principle provides explicit bounds on the adoption density at any time and location.
Verifiable prediction. For a technology diffusion process with power-law peer influence ($\alpha=2$) and time-varying advertising intensity $\mu(t)=\mu_{0}/(1+t)$ (diminishing returns), the total adoption level $\int_{\Omega}v(x,t)\,dx$ is bounded between $S(t)\int_{\Omega}u_{-}^{1/2}(x)\,dx$ and $S(t)\int_{\Omega}u_{+}^{1/2}(x)\,dx$, where $S(t)$ is given by (1.4). These bounds can be tested against empirical adoption curves for products such as smartphones, electric vehicles, or renewable energy technologies.
5.4. Engineering: groundwater flow and materials processing
In hydrology, the hydraulic head $v(x,t)$ in an unconfined aquifer with nonlinear permeability satisfies a porous medium equation of the form (1.1), where $\alpha$ depends on the soil type and $\mu(t)$ represents time-varying recharge or extraction rates [3]. The Bernoulli localization principle provides rigorous upper and lower bounds on the hydraulic head, which are useful for predicting well yields, assessing contamination risks, and designing groundwater management strategies.
In materials science, the sintering of powder compacts—a process by which loose particles consolidate into a dense solid under heat—involves density-dependent diffusion governed by equations of the form (1.1), where $v(x,t)$ represents the local density of the compact and $\mu(t)$ represents the time-dependent sintering temperature profile [12]. The localization bounds predict the achievable density range at any stage of the sintering process.
Verifiable prediction. For groundwater flow in a sandy aquifer ($\alpha\approx 2$) with seasonal recharge $\mu(t)=\mu_{0}(1+\cos(2\pi t))$, the hydraulic head at any monitoring well location $x$ is bounded by the explicit Bernoulli barriers. These bounds can be compared against field measurements from piezometric monitoring networks.
6. Image Denoising Application and Comparison with Perona–Malik
6.1. Motivation: from biological diffusion to image processing
From an applied perspective, one of the most striking consequences of the framework developed in this paper is its direct applicability to digital image processing. To our knowledge, this constitutes the first application of a porous-medium-type (Gurtin–MacCamy) diffusion equation to image denoising.
The connection between diffusion equations and image processing has been recognized since the seminal work of Perona and Malik [21, 22], who proposed anisotropic diffusion as a method for edge-preserving image smoothing. Their model, however, is based on a linear diffusion equation with a spatially varying, gradient-dependent diffusion coefficient. The Gurtin–MacCamy model, by contrast, employs a genuinely nonlinear power-law diffusion mechanism $\Delta v^{\alpha}$ that has fundamentally different mathematical properties: finite-speed propagation, degenerate parabolicity, and compact support. These properties make it naturally suited for image denoising, as the nonlinear diffusion automatically slows down near edges (where the intensity is low after appropriate preprocessing) and acts more aggressively in homogeneous regions.
6.2. Mathematical formulation for image denoising
We treat a grayscale image (or each color channel of a color image) as a function $v:\Omega_{h}\times[0,T]\rightarrow[0,1]$, where $\Omega_{h}\subset\mathbb{Z}^{2}$ is the discrete pixel grid. The denoising process is governed by the discrete version of the Gurtin–MacCamy equation:
$v_{i,j}^{n+1}=v_{i,j}^{n}+\Delta t\left[\Delta_{h}(v^{\alpha})_{i,j}-\mu(n\Delta t)\,v_{i,j}^{n}\right],$ (6.1)
where $\Delta_{h}$ denotes the standard five-point discrete Laplacian
$\Delta_{h}(u)_{i,j}=u_{i+1,j}+u_{i-1,j}+u_{i,j+1}+u_{i,j-1}-4u_{i,j}.$ (6.2)
In the denoising application, the term $\mu(t)\,v$ acts as a time-dependent damping that suppresses noise more aggressively at early stages (when the noise level is high) and less aggressively as the image converges to a clean state. The function $\mu(t)=0.3/(1+t)$ implements this strategy: strong initial absorption that decays over time.
6.3. Comparison with Perona–Malik anisotropic diffusion
The Perona–Malik (PM) model [21, 22] is the standard benchmark for edge-preserving image smoothing. It replaces the linear diffusion $\Delta v$ with the anisotropic operator
$\nabla\cdot\bigl(g(|\nabla v|)\,\nabla v\bigr),$ (6.3)
where $g(s)=\exp(-s^{2}/K^{2})$ is an edge-stopping function. The PM model reduces diffusion near edges (where $|\nabla v|$ is large) and maintains full diffusion in flat regions.
The key difference is that the GM model achieves edge preservation through a fundamentally different mechanism: the power-law nonlinearity $\Delta v^{\alpha}$ with $\alpha>1$ automatically reduces the diffusion coefficient when the intensity $v$ is small, without requiring explicit gradient estimation. This makes the GM model more robust to noise in the gradient computation, which is a well-known weakness of the PM approach at high noise levels.
6.4. Algorithmic description
The complete denoising pipeline is summarized in Algorithm 1.
Algorithm 1 Gurtin–MacCamy Image Denoising
0:  Clean image $I\in[0,1]^{M\times N\times 3}$, noise level $\sigma$, exponent $\alpha$, time step $\Delta t$, number of steps $K$, growth function $\mu(\cdot)$
0:  Denoised image $I_{\mathrm{den}}\in[0,1]^{M\times N\times 3}$
1:  Generate noisy image: $I_{\mathrm{noisy}}\leftarrow\mathrm{clip}(I+\sigma\cdot\mathcal{N}(0,1),\;0,\;1)$
2:  for each color channel $c\in\{R,G,B\}$ do
3:   $v\leftarrow I_{\mathrm{noisy}}[:,:,c]$
4:   for $n=0,1,\ldots,K-1$ do
5:    $v_{\alpha}\leftarrow v^{\alpha}$ (elementwise power)
6:    $L\leftarrow\Delta_{h}(v_{\alpha})$ (discrete Laplacian via (6.2))
7:    $v\leftarrow v+\Delta t\cdot(L-\mu(n\Delta t)\cdot v)$
8:    $v\leftarrow\mathrm{clip}(v,\;0,\;1)$
9:   end for
10:   $I_{\mathrm{den}}[:,:,c]\leftarrow v$
11:  end for
12:  Apply postprocessing (contrast stretch + gamma correction) to $I_{\mathrm{den}}$
13:  return $I_{\mathrm{den}}$
6.5. Python implementation: transparent explanation
We now present the complete Python implementation with a detailed explanation of each component. The goal is full transparency: every design choice is justified mathematically or algorithmically, so that the reader can reproduce, modify, and extend the code.
6.5.1 Imports and parameters
1import numpy as np
2import imageio.v2 as imageio
3import matplotlib.pyplot as plt
4from skimage.metrics import structural_similarity as ssim
5
6alpha_gm = 4
7dt_gm = 0.0118
8steps_gm = 400
9
10dt_pm = 0.2
11steps_pm = 40
12K_pm = 0.1
13
14gamma_correction = 1.05
15stretch_strength = 0.9
16brightness_boost = 1.0
17
18sigma = 0.18
19
20def mu(t):
21 return 0.3 / (1 + t)
Listing 1: Imports and model parameters.
Explanation. The code imports four libraries: numpy for array operations, imageio for reading image files, matplotlib for visualization, and skimage.metrics for the SSIM quality metric.
The Gurtin–MacCamy parameters are: $\alpha=4$ (the diffusion exponent, corresponding to the radiation-dominated regime), $\Delta t=0.0118$ (the Euler time step, chosen small enough for stability of the explicit scheme), and $K=400$ (the number of diffusion steps). The Perona–Malik parameters are: $\Delta t=0.2$, $K=40$, and $K_{\mathrm{pm}}=0.1$ (the edge-stopping threshold). The postprocessing parameters control contrast recovery after denoising (see below).
The noise level $\sigma=0.18$ specifies the standard deviation of the additive Gaussian noise. The function $\mu(t)=0.3/(1+t)$ implements the time-dependent growth rate from the theoretical framework: it provides strong damping at early times (when noise is dominant) and progressively weakens as the image approaches a clean state.
6.5.2 Image quality metrics
1def compute_mse(a, b):
2 return np.mean((a - b)**2)
3
4def compute_psnr(a, b):
5 mse = compute_mse(a, b)
6 if mse == 0:
7 return float("inf")
8 return 20 * np.log10(1.0 / np.sqrt(mse))
9
10def compute_ssim_color(a, b):
11 return np.mean([
12 ssim(a[:,:,c], b[:,:,c], data_range=1.0)
13 for c in range(3)
14 ])
Listing 2: Quality metrics: MSE, PSNR, and SSIM.
Explanation. Three standard image quality metrics are implemented.
  • Mean Squared Error (MSE): $\mathrm{MSE}=\frac{1}{MN}\sum_{i,j}(a_{i,j}-b_{i,j})^{2}$, measuring the average squared pixel difference. Lower values indicate better reconstruction.
  • Peak Signal-to-Noise Ratio (PSNR): $\mathrm{PSNR}=20\log_{10}(1/\sqrt{\mathrm{MSE}})$ dB, measuring the ratio of peak signal to noise on a logarithmic scale. Higher values indicate better quality; a difference of 3 dB corresponds to a factor-of-two improvement in MSE.
  • Structural Similarity Index (SSIM): A perceptual metric that compares luminance, contrast, and structural information between two images. Values range from 1-1 to $1$, with $1$ indicating perfect similarity. For color images, the SSIM is computed per channel and averaged.
6.5.3 Postprocessing
1def postprocess(u):
2 v = u.copy()
3 v_min = v.min()
4 v_max = v.max()
5 if v_max > v_min:
6 v = (v - v_min) / (v_max - v_min)
7 v = stretch_strength * v + (1 - stretch_strength) * u
8 v = np.power(v, gamma_correction)
9 v *= brightness_boost
10 return np.clip(v, 0, 1)
Listing 3: Postprocessing: contrast stretching and gamma correction.
Explanation. After the diffusion process, the image intensities may have shifted and compressed. The postprocessing restores visual quality through three operations:
@@ 6.5.3 Postprocessing
    Contrast stretching: The image is linearly mapped to the full $[0,1]$ range by $v\mapsto(v-v_{\min})/(v_{\max}-v_{\min})$. This compensates for the intensity reduction caused by the damping term $\mu(t)\,v$ in the diffusion equation.
@@
    Blending: The stretched image is blended with the original diffused image using the parameter stretch_strength=0.9=0.9:
    $v_{\mathrm{out}}=0.9\cdot v_{\mathrm{stretched}}+0.1\cdot v_{\mathrm{diffused}}.$
    This prevents over-enhancement of low-contrast regions.
@@
    Gamma correction: A mild gamma correction ($\gamma=1.05$) is applied to adjust the luminance curve, slightly darkening the midtones for a more natural appearance.
6.5.4 The Gurtin–MacCamy diffusion kernel
1def diffuse_channel_gm(u, alpha, dt, steps):
2 u = u.copy()
3 for n in range(steps):
4 u_alpha = u**alpha
5 lap = (
6 np.roll(u_alpha, 1, axis=0) +
7 np.roll(u_alpha, -1, axis=0) +
8 np.roll(u_alpha, 1, axis=1) +
9 np.roll(u_alpha, -1, axis=1) -
10 4*u_alpha
11 )
12 u = u + dt * (lap - mu(n*dt)*u)
13 u = np.clip(u, 0, 1)
14 return u
Listing 4: Gurtin–MacCamy nonlinear diffusion for a single color channel.
Explanation. This function implements the core of the Gurtin–MacCamy denoising algorithm, corresponding to the discrete evolution equation (6.1). At each time step $n$:
  1. 1.
    Compute $v^{\alpha}$: The elementwise power $u^{\alpha}$ is computed. For $\alpha=4$, this strongly attenuates low intensities (noise) while preserving high intensities (signal).
  2. 2.
    Compute the discrete Laplacian: The five-point stencil (6.2) is implemented via np.roll, which shifts the array along each axis. The sum of four shifted copies minus $4$ times the original gives the standard finite-difference approximation to $\Delta(v^{\alpha})$. The use of np.roll implements periodic boundary conditions, which is standard practice in image processing and avoids boundary artifacts.
  3. 3.
    Euler update: The forward Euler step
    $v^{n+1}=v^{n}+\Delta t\bigl[\Delta_{h}(v^{\alpha})-\mu(n\Delta t)\,v^{n}\bigr]$
    advances the solution by one time step. The damping term $\mu(n\Delta t)\,v^{n}$ implements the time-dependent absorption from the Gurtin–MacCamy framework.
@@ 6.5.4 The Gurtin–MacCamy diffusion kernel
    Clip to $[0,1]$: Pixel values are clamped to the valid range to maintain physical meaningfulness and numerical stability.
The mathematical connection to the theoretical framework is direct: this function implements a forward Euler discretization of the PDE $\partial_{t}v=\Delta v^{\alpha}-\mu(t)v$, which corresponds to equation (1.1) with the sign of the reaction term adapted for the absorption regime appropriate to denoising.
6.5.5 The Perona–Malik diffusion kernel
1def perona_malik_channel(u, dt, steps, K):
2 u = u.copy()
3 for _ in range(steps):
4 ux_f = np.roll(u, -1, axis=1) - u
5 ux_b = u - np.roll(u, 1, axis=1)
6 uy_f = np.roll(u, -1, axis=0) - u
7 uy_b = u - np.roll(u, 1, axis=0)
8
9 cxf = np.exp(-(ux_f**2) / (K**2))
10 cxb = np.exp(-(ux_b**2) / (K**2))
11 cyf = np.exp(-(uy_f**2) / (K**2))
12 cyb = np.exp(-(uy_b**2) / (K**2))
13
14 div = (cxf*ux_f - cxb*ux_b + cyf*uy_f - cyb*uy_b)
15 u = u + dt * div
16 u = np.clip(u, 0, 1)
17 return u
Listing 5: Perona–Malik anisotropic diffusion for a single color channel.
Explanation. This function implements the classical Perona–Malik anisotropic diffusion model [21, 22] for comparison. At each time step:
  1. 1.
    Compute finite differences: Forward and backward differences are computed along both spatial axes (horizontal: $x$, vertical: $y$), giving four gradient estimates per pixel.
  2. 2.
    Edge-stopping function: The Perona–Malik conductivity $g(s)=\exp(-s^{2}/K^{2})$ is applied to each gradient estimate. When the gradient is large (near an edge), $g\approx 0$ and diffusion is suppressed; when the gradient is small (in a flat region), $g\approx 1$ and full diffusion occurs.
  3. 3.
    Divergence and update: The discrete divergence of the flux $g\cdot\nabla v$ is computed and the Euler step is applied, as in the continuous equation $\partial_{t}v=\nabla\cdot(g(|\nabla v|)\nabla v)$.
The Perona–Malik model requires choosing the threshold $K$ carefully: too small, and noise is preserved; too large, and edges are blurred. This sensitivity is a recognized limitation of the PM approach, particularly at high noise levels where gradient estimates are unreliable.
6.5.6 The complete denoising pipeline
1img = imageio.imread("lenna.png").astype(np.float32) / 255.0
2gauss = sigma * np.random.randn(*img.shape)
3noisy_img = np.clip(img + gauss, 0, 1)
4
5den_gm = np.zeros_like(img)
6den_pm = np.zeros_like(img)
7
8for c in range(3):
9 den_gm[:,:,c] = diffuse_channel_gm(
10 noisy_img[:,:,c], alpha_gm, dt_gm, steps_gm
11 )
12 den_pm[:,:,c] = perona_malik_channel(
13 noisy_img[:,:,c], dt_pm, steps_pm, K_pm
14 )
15
16den_gm_pp = postprocess(den_gm)
17den_pm_pp = postprocess(den_pm)
18
19for label, im in [("Noisy", noisy_img),
20 ("Gurtin-MacCamy", den_gm_pp),
21 ("Perona-Malik", den_pm_pp)]:
22 print(f"{label}: MSE={compute_mse(img, im):.6f}, "
23 f"PSNR={compute_psnr(img, im):.2f} dB, "
24 f"SSIM={compute_ssim_color(img, im):.4f}")
Listing 6: Main pipeline: noise addition, denoising, and metric computation.
Explanation. The main pipeline operates in four stages:
  1. 1.
    Image loading and normalization: The test image (lenna.png) is loaded and normalized to $[0,1]$ by dividing by $255$.
  2. 2.
    Noise addition: Additive white Gaussian noise with standard deviation $\sigma$ is generated and added to the clean image. The result is clipped to $[0,1]$ to maintain valid pixel values.
  3. 3.
    Channel-wise denoising: Both the GM and PM denoising algorithms are applied independently to each of the three color channels (R, G, B). This channel-independent approach is standard in color image processing and avoids introducing color artifacts.
  4. 4.
    Postprocessing and evaluation: The postprocessing function restores contrast to the denoised images, and the three quality metrics (MSE, PSNR, SSIM) are computed for each method by comparing against the original clean image.
6.6. Results and comparison